OpenStudy (anonymous):

how to solve (m+8)(m-2)=0

OpenStudy (anonymous):

\[(m+8)(m-2)=0 \] explain how to please

OpenStudy (anonymous):

One of those terms equals zero, because no non-zero numbers multiply to give zero. The solutions are given by setting each term equal to zero and solving the resulting simple equations.

OpenStudy (anonymous):

I don't really remember F.O.I.L but I know that I can use it

OpenStudy (anonymous):

You don't need FOIL, just set each term equal to zero and solve for m.

OpenStudy (anonymous):

if I set each term to equal zero, I get two answers, do I just combine these answers so I'd get -6 ?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

No, each answer is a separate solution.

OpenStudy (anonymous):

solve for m?

OpenStudy (e.cociuba):

m= -8 m= 2

OpenStudy (e.cociuba):

0 = m + 8 Solve. 0= m - 2 Solve.

OpenStudy (anonymous):

\[(m+8)= 0 \] \[m = -8\] ------------------- \[(m-2)=0\] \[m = 2\]

OpenStudy (anonymous):

alright

OpenStudy (e.cociuba):

You got it? :)

OpenStudy (anonymous):

yes

OpenStudy (e.cociuba):

mm kay :))

OpenStudy (anonymous):

can you also help me solve another one?

OpenStudy (e.cociuba):

Ya, post a new question though :))

OpenStudy (anonymous):

Please post separate questions. It saves scrolling and allows people to receive separate medals for both answers.

OpenStudy (anonymous):

ok

OpenStudy (e.cociuba):

Yuppers @Numb3r1 :))

OpenStudy (e.cociuba):

@isaacme What the heck? :P

OpenStudy (anonymous):

lol ikr

OpenStudy (anonymous):

e.cciuba i use this rule deta=b²-4ac for solve kind of this equation ax+by+c =0

OpenStudy (e.cociuba):

The tops correct, but wats w/ the last part?

OpenStudy (anonymous):

lol no i use two ways.. just give him how to solve when be this formula ax²+by+c and will get the same answer. no faut : deta=b²-4ac=sqrt{100} \[m=\frac{ -6+\sqrt{100} }{ 2}=2 and m2=\frac{ -6-\sqrt{100} }{ 2}=8\]

OpenStudy (e.cociuba):

mmm kay! :))